Status Report

NOAA SATOPS Morning Report: Tuesday, May 22, 2012

By SpaceRef Editor
May 22, 2012
Filed under ,

SATOPS Morning Report: Tuesday, May 22, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 143

GOES
GOES-12 DOY142 intermittent Imager instrument cycle slips from through 10:00z to 14:58z.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
N-18 DOY143 06:56z G4A and 4B rev. 36090 from Wallops CDA were lost due to an incorrect ground station frequency setting for data recovery. Under investigation.

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
N-18 DOY143 06:56z G4A and 4B were lost. See POES notes above for more details.

Questions or comments to [email protected]

SpaceRef staff editor.